### v1

# SAGE: Sequence-level Adaptive Gradient Evolution for Generative Recommendation

Abstract.
While works such as OneRec have validated the scaling laws of Large Language Models (LLMs) in recommender systems, they rely on a cumbersome separate vocabulary. This dependency prevents the model architecture from reusing native LLM vocabularies, resulting in high maintenance costs and poor scalability. In response, we aim to efficiently reuse open-source LLM architectures without constructing a separate tokenization vocabulary. Furthermore, we identify that the optimization strategy of OneRec—Gradient-Bounded Policy Optimization (GBPO)—suffers from a ”Symmetric Conservatism” problem: its static gradient boundaries structurally suppress the update momentum required for cold-start items and fail to prevent diversity collapse in high-noise environments.To address this issue, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: By combining a geometric mean importance ratio with decoupled multi-objective advantages, we eliminate token-level variance and resolve the ”Reward Collapse” problem. (2) Asymmetric Adaptive Dynamics: We construct a dynamic gradient manifold that applies a ”Boost Factor” to high-potential cold-start items to achieve super-linear updates and employs an ”Entropy-Aware Penalty” to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
Generative Recommender Systems, Reinforcement Learning
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
In the digital age, recommender systems have become indispensable tools for user navigation of vast information landscapes. However, traditional methodologies, primarily leveraging historical interaction data and techniques such as collaborative filtering and matrix factorization, exhibit inherent shortcomings. Although effective in data-rich environments for capturing certain facets of user preference, these approaches often overlook valuable external knowledge, thereby contributing to the well-established ”filter bubble” effect (Koren et al., 2009). Moreover, industrial deployments frequently maintain a separation between model optimization and strategic iteration. This division, often manifested through distinct teams dedicated to model tuning versus strategy adjustment (Covington et al., 2016; Liu et al., 2017; Qin et al., 2022; Luo et al., 2024), can impede the rapid adaptation necessary to address evolving business requirements.
Recent advancements in large-scale pre-trained language models (LLMs) have introduced a transformative opportunity for the recommender system domain. These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data  (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025). Capitalizing on this potential to reshape data construction and model architecture in recommender systems holds the promise of overcoming the inherent limitations of traditional approaches, leading to enhancements in both the accuracy and diversity of personalized recommendations.
In generative recommendation systems such as OneRec, adapting an LLM to the recommendation task typically requires constructing a dedicated vocabulary to convert videos into Semantic IDs. Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges.
First, semantic collision is a major issue: multiple distinct videos may be mapped to the exact same token sequence, while a large portion of tokens remain unmapped. This can lead to collision rates exceeding 30%, meaning that the model is unable to distinguish 30% of the items, resulting in severe information loss  (Zhou et al., 2025a). Second, in industrial-scale scenarios where hundreds of millions of new videos are uploaded daily, maintaining the tokenizer incurs substantial operational overhead.
Moreover, LLM adaptation training is mandatory. A dedicated training stage is required for the LLM to understand item tokens and absorb collaborative signals. General-purpose LLMs cannot interpret these newly defined “Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a). This includes initializing embeddings for the new tokens and freezing other parameters during training so that the LLM can learn the semantics of video tokens.
To circumvent these issues, we explore text-based models that do not require constructing a separate tokenization vocabulary, enabling efficient reuse of open-source LLM architectures. The approach is outlined as follows:
Data Construction Optimized for Large Language Models.
The core advantage of this paradigm lies in the fact that it can directly adopt the architecture of large language models without the need to construct a separate tokenization vocabulary, thus enabling the most efficient reuse of open-source large language models. User-side: We integrate user profiles with historical behavior information, constructing prompt inputs that combine static user attributes and dynamic interaction data. This ensures that the large model can capture both the basic characteristics and behavioral habits of the user. Item-side: We comprehensively leverage item descriptions, titles, and content-derived tags to provide a multi-faceted semantic representation of the items. Furthermore, we generate positive samples by constructing an item sequence based on the user’s subsequent interactions, guided by these tags. Label Construction: We generate the training label, an item sequence that accurately reflects the user’s preferences, based on their subsequent behavior data. This provides a more precise supervisory signal for model training.
Two-Stage Interactive Training Strategy.
SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of DeepSeek R1 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions. This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation.Additionally, high-quality samples generated in the RLHF stage will be added to the training samples. RLHF Stage: Building upon the SFT stage, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: We propose a geometric mean importance ratio combined with a decoupled multi-objective advantage estimator. This eliminates token-level variance and resolves ”Reward Collapse” in multi-objective scenarios (e.g., balancing clicks and duration). (2) Asymmetric Adaptive Dynamics: SAGE constructs a dynamic gradient manifold that is asymmetric by design. It introduces a ”Boost Factor” to allow super-linear updates for high-potential cold-start items, and an ”Entropy-Aware Penalty” that dynamically intensifies gradients for low-diversity negative samples to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
Figure 1. Illustrates the motivation behind our research
Overall, Figure 1 illustrates the motivation behind our research. Our approach offers a novel perspective on recommender system design by seamlessly integrating the strengths of large language models and reinforcement learning. It effectively addresses the limitations of traditional methods—particularly the issue of information silos—while providing a solid technical foundation for multi-objective and customized recommendation in real-world business scenarios.
2. Related Work
2.1. Generative Recommendation and Its Challenges
With the breakthroughs in generative artificial intelligence, the paradigm of recommender systems is shifting from traditional Discriminative Ranking to Generative Recommendation. The core idea of generative recommendation is to reformulate the task as an end-to-end autoregressive generation problem, directly generating the Semantic ID sequence of the target item, thereby enabling direct optimization toward the final objective. This requires a tokenization strategy that maps items into discrete token sequences (Semantic IDs). OneRec is a pioneering paradigm in this direction, relying on predefined structured vocabularies  (Li et al., 2025; Deng et al., 2025).
However, this paradigm introduces a dependency on a heavy vocabulary: the model must memorize the mapping between semantic tokens and specific items. The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b).
2.2. Evolution of Policy Optimization
To align large models with complex business objectives, reinforcement learning–based methods have become mainstream. GRPO  (Shao et al., 2024) eliminates the need for a Critic network by normalizing rewards within sampled groups, and it has been widely applied in reasoning tasks. However, GRPO suffers from the problem of Reward Collapse in multi-objective optimization, where samples with different performance may end up with identical advantage values after normalization, reducing the resolution of training signals. In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b).
To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization). GBPO introduces a dynamic gradient boundary derived from binary cross-entropy (BCE), effectively preventing gradient explosion caused by negative samples and improving stability in large-scale training.
Although GBPO resolves the stability issue, its symmetric and conservative design ($r\leq 1$) introduces new bottlenecks. Cold-start video views dropped by 44.7%, because the strict upper bound on gradients prevents strong updates on sparse positive signals for new items. Cluster density increased by 11.7%, as the static boundary fails to preserve entropy-increasing exploratory behavior.
Therefore, we propose SAGE, a unified optimizer that is sequence-level, adaptive, and multi-objective.Token-level optimization introduces noise in Semantic ID generation. Following the logic of GSPO  (Zheng et al., 2025), SAGE elevates the importance ratio to the sequence level (geometric mean), ensuring the overall consistency of tokenized item IDs. BAPO  (Xi et al., 2025) shows that standard clipping suppresses exploration. SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-V2. To offset GBPO’s suppression of new items, SAGE incorporates the Clip-Higher philosophy from DAPO  (Yu et al., 2025), enabling super-linear updates for positive cold-start signals. To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and watch time).
3. Method
The proposed recommender system framework comprises three primary components: Data Construction, the SFT Stage, and the RLHF Stage, as depicted in Figure  2.
Figure 2. An overview of the modeling approach used in RecLLM-R1.
3.1. Data Construction
To align with the input requirements of large language models, we constructed a training dataset specifically tailored for recommender systems by transforming features into natural language descriptions. User-side Data:We transformed user profiles (e.g., age, gender, interests) and recent interaction history (e.g., sequences of clicked, viewed, liked, bookmarked, or purchased items) into chronologically ordered textual descriptions, aggregated by topic. For instance, ”Over the past 7 days, the user viewed ’Sichuan Cuisine’ videos twice and bookmarked one article titled ’Chengdu Travel Guide’.” Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions. Label Construction: Based on user engagement metrics such as browsing duration and interaction behaviors (e.g., bookmarking, sharing, commenting, following, liking), we constructed a sequence of items that reflects the user’s eventual preferences as positive training samples. This ensures the authenticity and business relevance of the labels. Prompt Construction: We concatenated the user-side and item-side textual information. The objective is to explicitly recommend the most relevant sequence of items to the user, with candidate items drawn from the provided data pool.
Through this methodology, we transformed traditional structured and semi-structured recommendation data into a natural language format interpretable and processable by large language models (LLMs). This data construction approach not only preserves the benefits of traditional interaction data but also imbues large language models with rich semantic context, thereby mitigating the limitations of the ”filter bubble” effect.
3.2. SFT Stage: Activating the Recommender Capability of the Large Language Model
Following the construction of the textualized recommendation dataset, we proceeded to the Supervised Fine-Tuning (SFT) stage. The primary objective of this phase was not to immediately achieve optimal recommendation performance, but rather to ”activate” or ”cold-start” the pre-trained large language model (LLM), endowing it with the foundational ability to address recommendation tasks.
For efficiency, we used the pre-trained Qwen3-8B model and adopted the previously constructed dataset that is compact in scale yet high in quality, which includes high-quality sample data generated in the RLHF stage. In this stage, the emphasis was on leveraging a minimal amount of data to rapidly unlock the LLM’s potential within the recommendation domain, enabling it to learn the requisite input-output formats and develop a rudimentary understanding of recommendation task modeling. This process prepares the model for the subsequent, more intricate multi-step optimization.
Essentially, the SFT stage serves as a ”warm-up” providing the model with baseline recommendation capabilities prior to reinforcement learning. While SFT can already address certain recommendation needs—particularly cold-start scenarios, owing to the LLM’s general knowledge—it is inherently insufficient for optimizing complex long-term objectives, balancing multiple recommendation facets (such as diversity and novelty), or executing advanced reasoning. Therefore, SFT acts as a critical precursor, priming the model for comprehensive optimization in the subsequent RLHF stage. This design closely mirrors the training paradigm of DeepSeek R1.
3.3. RLHF Stage: Preference Alignment with Real-World User Interactions
After establishing foundational capabilities in the Supervised Fine-Tuning (SFT) stage, we introduce the RLHF phase of model. To address the cold-start and diversity issues mentioned in OneRec-V2, we propose SAGE (Sequence-level Adaptive Gradient Evolution). The core contribution of SAGE is the Asymmetric Dynamic Bounding mechanism, which adjusts the optimization trajectory based on the nature of the feedback (Positive vs. Negative) and the entropy of the generated list.
3.3.1. Adaptive Gradient Evolution
SAGE Objective Function. To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term ($1/L$):
(1) $\begin{split}\nabla\mathcal{J}_{SAGE}(\theta)&=\mathbb{E}_{\tau\sim\pi_{old}}\Bigg[\frac{r_{slate}(\theta)}{\Phi(S,A)}\cdot A_{SAGE}\\ &\cdot\underbrace{\frac{1}{L}\sum_{t=1}^{L}\nabla\log\pi_{\theta}(i_{t}|u,S_{<t})}_{\text{Item-Averaged Gradient}}\Bigg]\end{split}$
  • $\frac{1}{L}$: Length Normalization Factor. This ensures the gradient magnitude aligns with the scale of $r_{slate}$ ($\approx 1$), preventing the optimizer from being biased towards longer sequences (though $L$ is fixed here, this maintains mathematical rigor).
  • $\Phi(S,A)$: The Adaptive Dynamic Denominator, defined below.
Case A: Positive Boost for Cold-Start (Solving Suppression). When the slate receives positive feedback ($A_{SAGE}\geq 0$), standard GBPO restricts the update ratio to $r\leq 1$. This is detrimental for cold-start items where the initial probability $\pi_{old}$ is extremely low. SAGE introduces a Boost Factor:
(2) $\Phi_{pos}=\begin{cases}\frac{1}{1+\epsilon_{boost}}&\text{if }r_{slate}(\theta)>1+\epsilon_{boost}\\ 1&\text{otherwise}\end{cases}$
  • $\epsilon_{boost}$: The Positive Boost Factor (e.g., $0.2\sim 0.5$).
  • Mechanism: If the model successfully identifies a high-potential item (resulting in a high ratio $r_{slate}$), SAGE locks the denominator such that the effective gradient coefficient is maintained at $1+\epsilon_{boost}$.
  • Impact: This allows the probability of successful cold-start items to grow super-linearly, providing the necessary momentum to break through the initial exposure threshold.
Case B: Diversity-Aware Penalty (Solving Collapse). When the slate receives negative feedback ($A_{SAGE}<0$), we distinguish between ”failed exploration” and ”homogenized repetition.” We utilize the List Entropy $H(S)$ to measure the semantic diversity of the slate.
(3) $\Phi_{neg}=\max\left(1,\frac{1-r_{slate}(\theta)}{\lambda(H(S))}\right)$
The regulation factor $\lambda$ is negatively correlated with diversity:
(4) $\begin{split}\lambda(H(S))=1+\beta\cdot\\ \tanh\left(\max(0,\bar{H}-H(S))\right)\end{split}$
  • $H(S)$: The content diversity score of the current slate (e.g., category entropy).
  • $\bar{H}$: The moving average of historical diversity.
  • $\beta$: The Diversity Temperature.
  • Mechanism:
    • Cocoon Penalty ($H(S)<\bar{H}$) If the slate is homogenized (low entropy) and rejected, $\lambda>1$. This shrinks the denominator $\Phi_{neg}$, resulting in a larger gradient (stronger penalty). This forces the model to escape the local optimum of recommending safe but boring items.
    • Exploration Tolerance ($H(S)\geq\bar{H}$) If the slate is diverse (high entropy) but rejected, $\lambda\approx 1$. The boundary reverts to the standard GBPO form, treating the error as a standard training sample without excessive punishment.
We illustrate the Comparison of Gradient Boundaries in  fig. 3 and the Evolution of Training Dynamics in  fig. 4.
Figure 3. Comparison of Gradient Boundaries. This figure illustrates the effective gradient coefficient ($r_{slate}/\Phi$) as a function of the importance ratio ($r_{slate}$). GBPO (Dashed Red) applies a static symmetric bound, capping positive updates at 1.0. In contrast, SAGE (Solid Blue) employs an asymmetric dynamic bound: it boosts high-potential positive samples beyond 1.0 (Case A) and applies entropy-aware penalties to negative samples (Case B), imposing stronger penalties on low-diversity slates (light blue) compared to high-diversity ones (dark blue).
Figure 4. Evolution of Training Dynamics. (a) SAGE rapidly boosts the generation probability of cold-start items compared to GBPO, demonstrating the effectiveness of the Positive Boost mechanism. (b) SAGE maintains consistently higher list diversity throughout training, verifying that the Entropy-Aware Penalty successfully prevents mode collapse, whereas GBPO tends to converge towards lower-entropy states.
3.3.2. Multi-Objective Signal Decoupling
To construct a high-fidelity learning signal from complex user feedback (clicks, duration, comments), we adopt a Decouple-then-Aggregate strategy inspired by GDPO.
Decoupled Slate Advantage Directly summing weighted rewards leads to ”Reward Collapse,” where distinct behaviors map to identical advantage values. Instead, we normalize each objective independently within the sampling group before aggregation:
(5) $\begin{split}A_{SAGE}(S)=\text{Norm}_{batch}\Bigg(\sum_{m=1}^{M}w_{m}\\ \cdot\frac{R_{m}(S)-\mu_{m}^{group}}{\sigma_{m}^{group}+\epsilon}\Bigg)\end{split}$
  • $R_{m}(S)$: The raw reward value for the $m$-th objective (e.g., watch time).
  • $\mu_{m}^{group},\sigma_{m}^{group}$: The mean and standard deviation of rewards within the sampled group $G$ for a single user query.
  • $\text{Norm}_{batch}$: A final batch-level normalization ($z$-score) to ensure $A_{SAGE}$ follows a standard normal distribution, providing a stable scale for the gradient boundaries.
Sequence-level Ratio Formulation Since user feedback is directed at the item level, token-level importance ratios introduce high variance. We define the importance ratio as the Geometric Mean of the token probability ratios over the generated list:
(6) $\begin{split}r_{slate}(\theta)=\exp\Bigg(\frac{1}{L}\sum_{t=1}^{L}\log\\ \frac{\pi_{\theta}(i_{t}|u,S_{<t})}{\pi_{\theta_{old}}(i_{t}|u,S_{<t})}\Bigg)\end{split}$
  • $L$: The fixed length of the recommendation slate ($L=6$).
  • Physical Meaning: This metric smooths out intra-item token fluctuations, representing the average probability gain of generating the entire slate under the new policy versus the old policy.
4. Experiments
In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed framework. We aim to validate whether our Native-LLM + SAGE paradigm (”RecLLM”) can outperform state-of-the-art generative baselines in terms of accuracy, while simultaneously resolving the cold-start and diversity bottlenecks inherent in the OneRec-V2 paradigm.
4.1. Experimental Setup
4.1.1. Datasets
Public Datasets: We evaluated the proposed framework on three publicly available real-world benchmarks from the Amazon Product Reviews dataset  (He and McAuley, 2016), which encompasses user reviews and item metadata spanning from May 1996 to July 2014. Specifically, we utilized three categories of this dataset for the sequential recommendation task: “Sports” “Beauty” and “Toys”.
4.1.2. Baselines
We compare our method against distinct categories of baselines:
  • SASRec: The classic self-attentive sequential model.
  • Generative/Retrieval Models: ReaRec, TIGER, and LC-Rec, which utilize VQ-VAE or retrieval-augmented generation.
  • OneRec Variants:
    • OneRec-V2: The SOTA generative baseline using Semantic IDs and GBPO optimizer.
    • OneRec-SAGE: Applying our SAGE optimizer to the traditional OneRec (Semantic ID) architecture to test optimizer effectiveness in isolation.
    • Only SFT (RecLLM): The Native-LLM model trained only with Supervised Fine-Tuning.
    • RecLLM: The proposed framework combining the Native-LLM architecture with the SAGE optimizer.
4.1.3. Evaluation Metrics
We evaluate performance across three dimensions using top-$K$ metrics (default $K=10$):
  • Accuracy: Normalized Discounted Cumulative Gain (NDCG@K) and Recall (Recall@K).
  • Diversity: Intra-List Diversity (ILD), measuring the average semantic distance between items within a generated slate, and Entropy@K, measuring the distribution of recommended categories across the test set.Since all items in a dataset category belong to the same root (e.g., ”Beauty”), calculating entropy at the root level is trivial. Therefore, we utilize the second-level sub-categories (e.g., ”Skin Care”, ”Makeup”, ”Fragrance” for Beauty) provided in the item metadata to calculate Shannon Entropy. This measures whether the model recommends a diverse range of granular interests rather than converging to a single sub-niche.
  • Cold-Start: Given that precise item publication timestamps are unavailable in this dataset, we adopt the standard academic definition for cold-start evaluation. We define Cold Items as the set of items belonging to the bottom 20% of the interaction frequency distribution in the training set. This metric evaluates the model’s ability to discover and recommend long-tail items that have sparse historical interactions, preventing the ”rich-get-richer” feedback loop.
4.1.4. Implementation Details
The base model is Qwen3-8B. We set the slate length $L=6$. The SFT stage is trained for 1 epochs. In the RLHF stage, we use a learning rate of 5e65e-6 with the AdamW optimizer. For SAGE specific hyperparameters, based on validation set tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$, the Diversity Temperature $\beta=0.5$.
4.2. Main Results
Superiority of RecLLM: As shown in Table  1, our proposed method (”RecLLM”) consistently outperforms all baselines across all datasets. Specifically, compared to the strongest baseline OneRec-V2, RecLLM achieves improvements of 4.3% to 8.9% in NDCG metrics. This confirms that directly reusing the Native LLM vocabulary—when optimized correctly—can surpass models relying on heavy, separate tokenization.
The Necessity of SAGE: Notably, the ”Only SFT” model performs poorly, significantly lagging behind even SASRec. SFT mainly functions to activate the model’s fundamental capabilities. without reinforcement learning to align the Native LLM with recommendation goals, the model fails to capture complex user preferences. Furthermore, OneRec-SAGE (applying SAGE to the old architecture) shows mixed results: while competitive, it slightly underperforms OneRec-V2 in some metrics (e.g., R@5 on Beauty). This suggests that SAGE’s aggressive exploration requires the robust semantic representation of the RecLLM to fully realize its potential without destabilizing the training.
Table  2 investigates the impact of optimization strategies on recommendation diversity (Entropy) and long-tail item discovery (Cold-Recall).
Breaking the Symmetric Conservatism: Comparing RecLLM against OneRec-V2, we observe substantial gains:
  • Cold-Recall: SAGE improves cold-start performance by 49% (Beauty), 63% (Sports) and 43% (Toys). This validates that the Positive Boost Factor ($\epsilon_{boost}$) effectively allows high-potential fresh items to break through the symmetric gradient barrier of GBPO.
  • Diversity: Entropy improves by 5.9% , 10.36% and 7.36%, confirming that the Entropy-Aware Penalty prevents the model from collapsing into safe, popular items.
The Accuracy-Diversity Trade-off: It is worth noting that OneRec-SAGE achieves the highest absolute values for Entropy and Cold-Recall. However, revisiting Table  1, we see that OneRec-SAGE suffers a drop in top-k accuracy (e.g., lower R@5 than OneRec-V2). This indicates that applying aggressive exploration on the limited semantic ID vocabulary can lead to over-exploration. In contrast, RecLLM achieves the most Optimal state: it delivers the highest overall accuracy table 1 while still providing a massive upgrade in diversity and cold-start handling compared to the current SOTA baseline (OneRec-V2).
Table 1. Overall performance comparison on three Amazon Product Reviews datasets. Bold indicates the best performance, and underline indicates the second best.
Dataset Beauty Sports Toys
Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
SASRec 0.0393 0.0639 0.0209 0.289 0.0240 0.0389 0.0130 0.0178 0.0420 0.0658 0.0217 0.0294
ReaRec 0.0488 0.0702 0.0341 0.0231 0.0348 0.0152 0.0189 0.0249 0.0517 0.0706 0.0369 0.0430
TIGER 0.0413 0.0628 0.0277 0.0346 0.0216 0.0331 0.0145 0.0182 0.0367 0.0527 0.0255 0.0307
LC-Rec 0.0495 0.0764 0.0338 0.0424 0.0269 0.0418 0.0177 0.0225 0.0350 0.0549 0.0221 0.0285
OneRec-V2 0.0646 0.0929 0.0456 0.0550 0.0365 0.0551 0.0252 0.0317 0.0693 0.0958 0.0496 0.0581
OneRec-SAGE 0.0612 0.0931 0.0424 0.0557 0.0351 0.0541 0.0246 0.0297 0.0671 0.0960 0.0466 0.0583
Only SFT (RecLLM) 0.0291 0.023 0.0158 0.204 0.0163 0.0263 0.0104 0.0101 0.0386 0.0571 0.0169 0.0212
RecLLM 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
Improve +5.67% +4.7% +8.9% +7.2% +5.46% +4.33% +7.35% +6.96% +6.5% +6.3% +7.48% +7.11%
Table 2. Relative to OneRec‑v2, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
Dataset Beauty Sports Toys
Method Entropy@10 Cold-Recall Entropy@10 Cold-Recall Entropy@10 Cold-Recall
OneRec-V2 2.285 0.0326 2.274 0.0282 2.315 0.0365
OneRec-SAGE 2.551 0.0619 2.528 0.0713 2.574 0.0695
RecLLM 2.420 0.0487 2.510 0.0460 2.485 0.0522
Improve +5.9% +49% +10.36% +63% +7.36% +43%
4.3. Ablation Study
To understand the contribution of each component, we conduct an ablation study on the Amazon-Beauty dataset. The results are visually presented in Figure  5.
Figure 5. Ablation Study on Amazon-Beauty Dataset. Metrics are normalized to the full SAGE model. Removing individual components leads to degradation in their respective targeted areas (e.g., removing Boost hurts Cold-Start most), validating the design of each module.
  • w/o Positive Boost: Setting $\epsilon_{boost}=0$ reverts the positive boundary to standard GBPO ($r\leq 1$). We observe a drastic drop in Cold-Start Recall (Green Bar) to 0.50 of the full model’s performance, confirming this component’s critical role in item exploration.
  • w/o Entropy Penalty: Removing the entropy term and using the static GBPO boundary for negative feedback leads to a significant decrease in Entropy@10 (Orange Bar) to 0.78, indicating a collapse towards redundant recommendations.
  • w/o Signal Decoupling: Using a simple weighted sum of rewards instead of decoupled normalization results in a noticeable drop in NDCG@10 (Blue Bar) to 0.90 and a decrease in other metrics. This suggests that ”Reward Collapse” was indeed hindering the precise multi-objective optimization.
The full SAGE model achieves the best balance, demonstrating that the asymmetric adaptive dynamics and signal decoupling are complementary and essential.
5. Conclusion
In this paper, we identified fundamental limitations in existing optimization strategies for generative recommendation, specifically the ”Symmetric Conservatism” of Gradient-Bounded Policy Optimization (GBPO), which leads to cold-start suppression and diversity collapse. To address these challenges within a list-wise generation paradigm, we proposed SAGE (Sequence-level Adaptive Gradient Evolution).
SAGE introduces a novel theoretical framework featuring **Asymmetric Adaptive Dynamics**. By applying a ”Positive Boost” to high-potential successful slates and an ”Entropy-Aware Penalty” to low-diversity failures, SAGE dynamically aligns the optimization trajectory with the long-tailed, high-noise nature of recommendation feedback. Furthermore, we incorporated Multi-Objective Signal Decoupling to ensure high-fidelity gradient signals in complex real-world scenarios.
RecLLM demonstrates that native‑vocabulary reuse not only surpasses models dependent on large standalone tokenizers but also maximizes the utility of the underlying LLM. While the absence of an independent vocabulary necessitates language‑level item encoding—posing challenges for real‑time systems—this constraint is expected to fade with increasing computational capacity.

### v2

@@
 Abstract.
-While works such as OneRec have validated the scaling laws of Large Language Models (LLMs) in recommender systems, they rely on a cumbersome separate vocabulary.
-This dependency prevents the model architecture from reusing native LLM vocabularies, resulting in high maintenance costs and poor scalability.
-In response, we aim to efficiently reuse open-source LLM architectures without constructing a separate tokenization vocabulary.
-Furthermore, we identify that the optimization strategy of OneRec—Gradient-Bounded Policy Optimization (GBPO)—suffers from a ”Symmetric Conservatism” problem: its static gradient boundaries structurally suppress the update momentum required for cold-start items and fail to prevent diversity collapse in high-noise environments.To address this issue, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation.
-SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: By combining a geometric mean importance ratio with decoupled multi-objective advantages, we eliminate token-level variance and resolve the ”Reward Collapse” problem.
-(2) Asymmetric Adaptive Dynamics: We construct a dynamic gradient manifold that applies a ”Boost Factor” to high-potential cold-start items to achieve super-linear updates and employs an ”Entropy-Aware Penalty” to break information cocoons.
-Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
+Reinforcement learning–based preference optimization is increasingly used to align list-wise generative recommenders with complex, multi-objective user feedback, yet existing optimizers such as Gradient-Bounded Policy Optimization (GBPO) exhibit structural limitations in recommendation settings.
+We identify a Symmetric Conservatism failure mode in which symmetric update bounds suppress learning from rare positive signals (e.g., cold-start items), static negative-sample constraints fail to prevent diversity collapse under rejection-dominated feedback, and group-normalized multi-objective rewards lead to low-resolution training signals.
+To address these issues, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimizer designed for list-wise generative recommendation.
+SAGE introduces sequence-level signal alignment via a geometric-mean importance ratio and a decoupled multi-objective advantage estimator to reduce token-level variance and mitigate reward collapse, together with asymmetric adaptive bounding that applies positive Boost updates to successful slates and an entropy-aware penalty to discourage low-diversity failures.
+Experiments on Amazon Product Reviews and the large-scale RecIF-Bench demonstrate consistent improvements in top-K accuracy, cold-start recall, and diversity across both Semantic-ID and native-text action spaces, while preserving numerical stability during training.
+These results suggest that asymmetric, sequence-aware policy optimization provides a principled and effective framework for addressing optimization failures in generative recommendation.
 Generative Recommender Systems, Reinforcement Learning
-copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
 1. Introduction
@@
 Recent advancements in large-scale pre-trained language models (LLMs) have introduced a transformative opportunity for the recommender system domain.
-These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data  (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025).
+These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025; Wei et al., 2022; Kojima et al., 2022).
 Capitalizing on this potential to reshape data construction and model architecture in recommender systems holds the promise of overcoming the inherent limitations of traditional approaches, leading to enhancements in both the accuracy and diversity of personalized recommendations.
 In generative recommendation systems such as OneRec, adapting an LLM to the recommendation task typically requires constructing a dedicated vocabulary to convert videos into Semantic IDs.
-Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges.
+Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges  (Zhou et al., 2025b, a; Rajput et al., 2023).
 First, semantic collision is a major issue: multiple distinct videos may be mapped to the exact same token sequence, while a large portion of tokens remain unmapped.
@@
 A dedicated training stage is required for the LLM to understand item tokens and absorb collaborative signals.
-General-purpose LLMs cannot interpret these newly defined “Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a).
+General-purpose LLMs cannot interpret these newly defined ”Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a; Artetxe et al., 2020).
 This includes initializing embeddings for the new tokens and freezing other parameters during training so that the LLM can learn the semantics of video tokens.
-To circumvent these issues, we explore text-based models that do not require constructing a separate tokenization vocabulary, enabling efficient reuse of open-source LLM architectures.
+To evaluate whether the optimization issues are intrinsic to the optimizer rather than a specific item tokenization, we consider two representative action spaces in generative recommendation: (i) Semantic-ID generation as in OneRec, and (ii) native-text item rendering without a separate item vocabulary.
 The approach is outlined as follows:
-Data Construction Optimized for Large Language Models.
+System instantiation for evaluation: data construction.
 The core advantage of this paradigm lies in the fact that it can directly adopt the architecture of large language models without the need to construct a separate tokenization vocabulary, thus enabling the most efficient reuse of open-source large language models.
@@
 This provides a more precise supervisory signal for model training.
-Two-Stage Interactive Training Strategy.
-SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of DeepSeek R1 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions.
-This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation.Additionally, high-quality samples generated in the RLHF stage will be added to the training samples.
+System instantiation for evaluation: two-stage training.
+SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of Qwen3 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions.
+This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation.
+Additionally, high-quality samples generated in the RLHF stage will be added to the training samples.
 RLHF Stage: Building upon the SFT stage, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation.
@@
 Figure 1. Illustrates the motivation behind our research
+Conceptual illustration of the proposed approach and its motivation for combining large language models with reinforcement learning for recommendation.
+Contributions.
+We identify a practical failure mode of GBPO in generative recommendation, termed Symmetric Conservatism, where symmetric conservative bounds suppress rare-item positive learning signals and can trigger diversity collapse under rejection-dominated feedback.
+To address this, we propose SAGE, a sequence-level, multi-objective policy optimization framework for list-wise generative recommendation.
+SAGE combines a sequence-level importance ratio for variance reduction, an asymmetric adaptive bounding rule with a positive boost and an entropy-aware penalty, and a decoupled normalization scheme for multi-objective rewards.
+We evaluate SAGE on both academic and industrial benchmarks and observe consistent gains in accuracy, cold-start discovery, and diversity, while maintaining stable training.
 Overall, Figure 1 illustrates the motivation behind our research.
@@
 However, this paradigm introduces a dependency on a heavy vocabulary: the model must memorize the mapping between semantic tokens and specific items.
-The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b).
+The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b; Ranzato et al., 2015).
 2.2. Evolution of Policy Optimization
@@
 However, GRPO suffers from the problem of Reward Collapse in multi-objective optimization, where samples with different performance may end up with identical advantage values after normalization, reducing the resolution of training signals.
-In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b).
-To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization).
+In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b; Zhao et al., 2017; Afsar et al., 2022; Chen et al., 2019).
+To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization, OneRec-GBPO).
 GBPO introduces a dynamic gradient boundary derived from binary cross-entropy (BCE), effectively preventing gradient explosion caused by negative samples and improving stability in large-scale training.
@@
 BAPO  (Xi et al., 2025) shows that standard clipping suppresses exploration.
-SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-V2.
+SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-GBPO.
 To offset GBPO’s suppression of new items, SAGE incorporates the Clip-Higher philosophy from DAPO  (Yu et al., 2025), enabling super-linear updates for positive cold-start signals.
-To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and watch time).
+To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and Comment).
+Positioning of SAGE.
+Conceptually, SAGE belongs to the PPO/GBPO family of bounded policy optimization methods, but differs in two aspects that are crucial for recommendation: (i) SAGE performs sequence-level importance weighting to match list-wise feedback and reduce token-level variance; (ii) SAGE introduces an asymmetric, sample-adaptive bounding rule—boosting positive successful slates while penalizing low-entropy failures more strongly—to address cold-start suppression and diversity collapse.
+For multi-objective feedback, SAGE adopts a decouple-then-aggregate advantage estimator to avoid reward collapse after group normalization.
 3. Method
-The proposed recommender system framework comprises three primary components: Data Construction, the SFT Stage, and the RLHF Stage, as depicted in Figure  2.
-Figure 2. An overview of the modeling approach used in RecLLM-R1.
-3.1. Data Construction
+3.1. Problem Setup
+We study list-wise generative recommendation as a slate generation problem  (Ie et al., 2019).
+Given a user context $c$ (e.g., profile and recent behaviors), a policy $\pi_{\theta}$ generates a slate $S=(i_{1},\dots,i_{L})$ autoregressively.
+User feedback is provided at the slate level and can be multi-objective, denoted by $\mathbf{R}(S)=\{R_{m}(S)\}_{m=1}^{M}$.
+Our goal is to optimize the expected utility under mixed feedback while maintaining stable training under rejection-dominated data:
+(1) $\max_{\theta}\;\mathbb{E}_{S\sim\pi_{\theta}(\cdot|c)}\left[\sum_{m=1}^{M}w_{m}R_{m}(S)\right].$
+We instantiate the proposed formulation using a two-stage training pipeline (SFT followed by RLHF) on textualized recommendation data.
+The central algorithmic contribution is the RLHF optimizer, SAGE, which we describe in detail next and illustrate in Figure  2.
+Figure 2. An overview of the modeling approach used in TextRec.
+3.2. Implementation: Data Construction
 To align with the input requirements of large language models, we constructed a training dataset specifically tailored for recommender systems by transforming features into natural language descriptions.
 User-side Data: We transformed user profiles (e.g., age, gender, interests) and recent interaction history (e.g., sequences of clicked, viewed, liked, bookmarked, or purchased items) into chronologically ordered textual descriptions, aggregated by topic.
-For instance, ”Over the past 7 days, the user viewed ’Sichuan Cuisine’ videos twice and bookmarked one article titled ’Chengdu Travel Guide’.” Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions.
+For instance, ”Over the past 7 days, the user viewed ”Sichuan Cuisine” videos twice and bookmarked one article titled ”Chengdu Travel Guide”.
+Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions.
 Label Construction: Based on user engagement metrics such as browsing duration and interaction behaviors (e.g., bookmarking, sharing, commenting, following, liking), we constructed a sequence of items that reflects the user’s eventual preferences as positive training samples.
@@
 This data construction approach not only preserves the benefits of traditional interaction data but also imbues large language models with rich semantic context, thereby mitigating the limitations of the ”filter bubble” effect.
-3.2. SFT Stage: Activating the Recommender Capability of the Large Language Model
+3.3. Implementation: SFT Warm-up
 Following the construction of the textualized recommendation dataset, we proceeded to the Supervised Fine-Tuning (SFT) stage.
@@
 After establishing foundational capabilities in the Supervised Fine-Tuning (SFT) stage, we introduce the RLHF phase of model.
-To address the cold-start and diversity issues mentioned in OneRec-V2, we propose SAGE (Sequence-level Adaptive Gradient Evolution).
+To address the cold-start and diversity issues mentioned in OneRec-V2 (OneRec-GRPO), we propose SAGE (Sequence-level Adaptive Gradient Evolution).
 The core contribution of SAGE is the Asymmetric Dynamic Bounding mechanism, which adjusts the optimization trajectory based on the nature of the feedback (Positive vs.
 Negative) and the entropy of the generated list.
+3.4.1. Background: GBPO in OneRec-V2
+We briefly review Gradient-Bounded Policy Optimization (GBPO) proposed in the OneRec-V2 Technical Report (Zhou et al., 2025b).
+GBPO is a bounded policy optimization method tailored for recommendation settings with rejection-dominated feedback.
+It controls update magnitude by restricting the effective importance ratio to a conservative interval, which improves numerical stability compared to vanilla policy gradient.
+However, the symmetric and conservative design of GBPO can induce a failure mode that we term Symmetric Conservatism: (i) the upper bound suppresses learning from rare positive slates (e.g., cold-start items), and (ii) static constraints on negative updates may fail to prevent diversity collapse under repeated rejections.
+These limitations motivate our asymmetric and entropy-aware extension, SAGE.
 3.4.2. Adaptive Gradient Evolution
+SAGE as asymmetric, sample-adaptive clipping.
+Let $r_{\text{slate}}(\theta)$ be the sequence-level importance ratio (defined later). SAGE optimizes a PPO-style surrogate with an asymmetric and entropy-aware clipping rule:
+(2) $\mathcal{L}_{\text{SAGE}}(\theta)=\mathbb{E}\left[\mathrm{clip}_{\text{SAGE}}\!\left(r_{\text{slate}}(\theta),A_{\text{SAGE}}(S),H(S)\right)\cdot A_{\text{SAGE}}(S)\right],$
+where $\mathrm{clip}_{\text{SAGE}}(\cdot)$ boosts positive updates up to $1+\epsilon_{\text{boost}}$ and applies stronger penalties to low-entropy rejected slates.
 SAGE Objective Function.
-To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term ($1/L$):
+To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term $1/L$:
 (3) $\begin{split}\nabla\mathcal{J}_{SAGE}(\theta)&=\mathbb{E}_{\tau\sim\pi_{old}}\Bigg[\frac{r_{slate}(\theta)}{\Phi(S,A)}\cdot A_{SAGE}\\ &\cdot\underbrace{\frac{1}{L}\sum_{t=1}^{L}\nabla\log\pi_{\theta}(i_{t}|u,S_{<t})}_{\text{Item-Averaged Gradient}}\Bigg]\end{split}$
-  • $\frac{1}{L}$: Length Normalization Factor.
-This ensures the gradient magnitude aligns with the scale of $r_{slate}$ ($\approx 1$), preventing the optimizer from being biased towards longer sequences (though $L$ is fixed here, this maintains mathematical rigor).
-  • $\Phi(S,A)$: The Adaptive Dynamic Denominator, defined below.
-Case A: Positive Boost for Cold-Start (Solving Suppression).
-When the slate receives positive feedback ($A_{SAGE}\geq 0$), standard GBPO restricts the update ratio to $r\leq 1$.
-This is detrimental for cold-start items where the initial probability $\pi_{old}$ is extremely low.
-SAGE introduces a Boost Factor:
-(2) $\Phi_{pos}=\begin{cases}\frac{1}{1+\epsilon_{boost}}&\text{if }r_{slate}(\theta)>1+\epsilon_{boost}\\ 1&\text{otherwise}\end{cases}$
-  • $\epsilon_{boost}$: The Positive Boost Factor (e.g., $0.2\sim 0.5$).
-  • Mechanism: If the model successfully identifies a high-potential item (resulting in a high ratio $r_{slate}$), SAGE locks the denominator such that the effective gradient coefficient is maintained at $1+\epsilon_{boost}$.
-  • Impact: This allows the probability of successful cold-start items to grow super-linearly, providing the necessary momentum to break through the initial exposure threshold.
-Case B: Diversity-Aware Penalty (Solving Collapse).
-When the slate receives negative feedback ($A_{SAGE}<0$), we distinguish between ”failed exploration” and ”homogenized repetition.” We utilize the List Entropy $H(S)$ to measure the semantic diversity of the slate.
-(3) $\Phi_{neg}=\max\left(1,\frac{1-r_{slate}(\theta)}{\lambda(H(S))}\right)$
+Here, the factor $1/L$ normalizes the gradient magnitude to match the scale of $r_{\mathrm{slate}}(\theta)$ (typically around $1$), avoiding an implicit preference toward longer sequences.
+The term $\Phi(S,A)$ is an adaptive dynamic denominator defined below.
+Case A: Positive boost for cold-start (mitigating suppression).
+When the slate receives positive feedback ($A_{\mathrm{SAGE}}\geq 0$), standard GBPO effectively caps the update ratio at $r\leq 1$, which can suppress cold-start items whose initial probability under $\pi_{\mathrm{old}}$ is extremely small.
+We introduce a boost factor
+(4) $\Phi_{\mathrm{pos}}=\begin{cases}\frac{1}{1+\epsilon_{\mathrm{boost}}},&\text{if }r_{\mathrm{slate}}(\theta)>1+\epsilon_{\mathrm{boost}},\\ 1,&\text{otherwise}.\end{cases}$
+With $\epsilon_{\mathrm{boost}}\in[0.2,0.5]$, once $r_{\mathrm{slate}}(\theta)$ exceeds $1+\epsilon_{\mathrm{boost}}$, the denominator is locked so that the effective gradient coefficient is maintained at $1+\epsilon_{\mathrm{boost}}$.
+This enables successful cold-start items to gain probability in a super-linear manner, providing sufficient momentum to pass the initial exposure threshold.
+Case B: Diversity-aware penalty (mitigating collapse).
+When the slate receives negative feedback ($A_{\mathrm{SAGE}}<0$), we distinguish failed exploration from homogenized repetition by using the list entropy $H(S)$ as a proxy for semantic diversity:
+(5) $\Phi_{\mathrm{neg}}=\max\!\left(1,\;\frac{1-r_{\mathrm{slate}}(\theta)}{\lambda(H(S))}\right).$
 The regulation factor $\lambda$ is negatively correlated with diversity:
-(4) $\begin{split}\lambda(H(S))=1+\beta\cdot\\ \tanh\left(\max(0,\bar{H}-H(S))\right)\end{split}$
-  • $H(S)$: The content diversity score of the current slate (e.g., category entropy).
-  • $\bar{H}$: The moving average of historical diversity.
-  • $\beta$: The Diversity Temperature.
-  • Mechanism:
-    • Cocoon Penalty ($H(S)<\bar{H}$) If the slate is homogenized (low entropy) and rejected, $\lambda>1$.
-This shrinks the denominator $\Phi_{neg}$, resulting in a larger gradient (stronger penalty).
-This forces the model to escape the local optimum of recommending safe but boring items.
-    • Exploration Tolerance ($H(S)\geq\bar{H}$) If the slate is diverse (high entropy) but rejected, $\lambda\approx 1$.
-The boundary reverts to the standard GBPO form, treating the error as a standard training sample without excessive punishment.
+(6) $\lambda(H(S))=1+\beta\cdot\tanh\!\left(\max(0,\bar{H}-H(S))\right),$
+where $H(S)$ denotes the slate diversity score (e.g., category entropy), $\bar{H}$ is the moving average of historical diversity, and $\beta$ is the diversity temperature.
+If $H(S)<\bar{H}$, then $\lambda>1$, which reduces the denominator and yields a larger gradient magnitude, thereby imposing a stronger penalty on “cocooning” behavior.
+Conversely, if $H(S)\geq\bar{H}$, then $\lambda\approx 1$ and the update reverts to the standard GBPO boundary, treating the rejection as a regular training signal without excessive punishment.
 We illustrate the comparison of gradient boundaries in fig. 3 and the evolution of training dynamics in fig. 4.
@@
 (b) SAGE maintains consistently higher list diversity throughout training, verifying that the Entropy-Aware Penalty successfully prevents mode collapse, whereas GBPO tends to converge towards lower-entropy states.
+Algorithmic summary.
+The above design can be implemented as a PPO-style on-policy update with (i) a sequence-level importance ratio, (ii) decouple-then-aggregate multi-objective advantages, and (iii) an asymmetric, entropy-aware clipping rule.
+For completeness, we summarize the training procedure in Algorithm LABEL:alg:sage_main.
+Algorithm 1 SAGE: Sequence-level Adaptive Gradient Evolution
+0: Old policy $\pi_{\theta_{\text{old}}}$, trainable policy $\pi_{\theta}$, batch of contexts $\mathcal{B}$, slate length $L$, group size $K$, objectives $\{R_{m}\}_{m=1}^{M}$, boost $\epsilon_{\text{boost}}$, diversity temperature $\beta$, clip param $\epsilon_{\text{clip}}$
+0: Updated parameters $\theta$
+1: Initialize running diversity target $\bar{H}$
+2:for each training step do
+3:  // Step 1: Sampling & Evaluation
+4:  Sample batch $\mathcal{B}\sim\mathcal{D}$
+5:  Generate $K$ slates $\{S_{u,k}\}$ for each user $u\in\mathcal{B}$ using $\pi_{\theta_{\text{old}}}$
+6:  Compute Rewards $\{R_{m}(S_{u,k})\}$ and Diversity $H(S_{u,k})$
+7:  // Step 2: Signal Decoupling (Eq. 7)
+8:  for each user $u\in\mathcal{B}$ do
+9:   for each objective $m$ do
+10:    Normalize $R_{m}$ within group $K$: $\hat{R}_{u,k,m}\leftarrow\frac{R_{u,k,m}-\mu_{u,m}}{\sigma_{u,m}+\epsilon}$
+11:   end for
+12:   Aggregate: $\hat{A}_{u,k}\leftarrow\sum_{m=1}^{M}w_{m}\cdot\hat{R}_{u,k,m}$
+13:  end for
+14:  Batch Normalize $\hat{A}$ over entire batch $\mathcal{B}$ to get final $A_{u,k}$
+15:  // Step 3: Asymmetric Adaptive Optimization
+16:  for each sample $(u,k)$ in batch do
+17:   Compute Sequence Ratio (Geo-Mean):
+18:   $r_{u,k}\leftarrow\exp\left(\frac{1}{L}\sum_{t=1}^{L}\log\frac{\pi_{\theta}(i_{t}|c,S_{<t})}{\pi_{\theta_{\text{old}}}(i_{t}|c,S_{<t})}\right)$
+19:   if $A_{u,k}\geq 0$ then
+20:    Case A (Positive Boost):
+21:    Set lower bound: $\epsilon_{-}\leftarrow\epsilon_{\text{clip}}$
+22:    Set upper bound: $\epsilon_{+}\leftarrow\epsilon_{\text{boost}}$ {Allow super-linear update}
+23:   else
+24:    Case B (Diversity Penalty):
+25:    Calculate regulation: $\lambda\leftarrow 1+\beta\cdot\tanh(\max(0,\bar{H}-H(S_{u,k})))$
+26:    Set lower bound: $\epsilon_{-}\leftarrow\lambda-1$ {Stronger penalty if low entropy}
+27:    Set upper bound: $\epsilon_{+}\leftarrow\epsilon_{\text{clip}}$
+28:   end if
+29:   Compute PPO Surrogate:
+30:   $r_{\text{clip}}\leftarrow\text{clip}(r_{u,k},1-\epsilon_{-},1+\epsilon_{+})$
+31:   $\ell_{u,k}\leftarrow\min(r_{u,k}A_{u,k},r_{\text{clip}}A_{u,k})$
+32:  end for
+33:  $\mathcal{L}_{\text{SAGE}}\leftarrow-\frac{1}{|\mathcal{B}|\cdot K}\sum\ell_{u,k}$
+34:  Update $\theta$ by descending $\nabla_{\theta}\mathcal{L}_{\text{SAGE}}$
+35:  Update $\bar{H}$ via EMA
+36:end for
 3.4.3. Multi-Objective Signal Decoupling
 To construct a high-fidelity learning signal from complex user feedback (clicks, duration, comments), we adopt a Decouple-then-Aggregate strategy inspired by GDPO.
-Decoupled Slate Advantage Directly summing weighted rewards leads to ”Reward Collapse,” where distinct behaviors map to identical advantage values.
-Instead, we normalize each objective independently within the sampling group before aggregation:
+Decoupled Slate Advantage.
+Directly summing weighted rewards can cause Reward Collapse, where distinct behaviors map to identical advantage values.
+To avoid this, we normalize each objective independently within the sampling group and then aggregate them.
+For a slate $S$, the advantage is defined as
 (7) $\begin{split}A_{SAGE}(S)=\text{Norm}_{batch}\Bigg(\sum_{m=1}^{M}w_{m}\\ \cdot\frac{R_{m}(S)-\mu_{m}^{group}}{\sigma_{m}^{group}+\epsilon}\Bigg)\end{split}$
-  • $R_{m}(S)$: The raw reward value for the $m$-th objective (e.g., watch time).
-  • $\mu_{m}^{group},\sigma_{m}^{group}$: The mean and standard deviation of rewards within the sampled group $G$ for a single user query.
-  • $\text{Norm}_{batch}$: A final batch-level normalization ($z$-score) to ensure $A_{SAGE}$ follows a standard normal distribution, providing a stable scale for the gradient boundaries.
-Sequence-level Ratio Formulation Since user feedback is directed at the item level, token-level importance ratios introduce high variance.
-We define the importance ratio as the Geometric Mean of the token probability ratios over the generated list:
+where $R_{m}(S)$ denotes the raw reward of the $m$-th objective (e.g., Click or Comment), and $\mu_{m}^{group},\sigma_{m}^{group}$ are the mean and standard deviation computed within the sampled group $G$ for the same user query.
+The operator $\text{Norm}_{batch}$ applies a final batch-level $z$-score normalization so that $A_{SAGE}$ stays on a stable scale (approximately standard normal), which helps maintain well-behaved gradient magnitudes.
+Sequence-level Ratio Formulation.
+Since user feedback is directed at the item level, token-level importance ratios can introduce high variance.
+We therefore define the importance ratio at the slate level as the geometric mean of token probability ratios across the generated list:
 (8) $\begin{split}r_{slate}(\theta)=\exp\Bigg(\frac{1}{L}\sum_{t=1}^{L}\log\\ \frac{\pi_{\theta}(i_{t}|u,S_{<t})}{\pi_{\theta_{old}}(i_{t}|u,S_{<t})}\Bigg)\end{split}$
-  • $L$: The fixed length of the recommendation slate ($L=6$).
-  • Physical Meaning: This metric smooths out intra-item token fluctuations, representing the average probability gain of generating the entire slate under the new policy versus the old policy.
-4. Experiments
-In this section, we conduct comprehensive experiments to evaluate the effectiveness of the proposed framework.
-We aim to validate whether our Native-LLM + SAGE paradigm (”RecLLM”) can outperform state-of-the-art generative baselines in terms of accuracy, while simultaneously resolving the cold-start and diversity bottlenecks inherent in the OneRec-V2 paradigm.
-4.1. Experimental Setup
-4.1.1. Datasets
-Public Datasets: We evaluated the proposed framework on three publicly available real-world benchmarks from the Amazon Product Reviews dataset  (He and McAuley, 2016), which encompasses user reviews and item metadata spanning from May 1996 to July 2014.
-Specifically, we utilized three categories of this dataset for the sequential recommendation task: “Sports” “Beauty” and “Toys”.
-4.1.2. Baselines
-We compare our method against distinct categories of baselines:
-  • SASRec: The classic self-attentive sequential model.
-  • Generative/Retrieval Models: ReaRec, TIGER, and LC-Rec, which utilize VQ-VAE or retrieval-augmented generation.
-  • OneRec Variants:
-    • OneRec-V2: The SOTA generative baseline using Semantic IDs and GBPO optimizer.
-    • OneRec-SAGE: Applying our SAGE optimizer to the traditional OneRec (Semantic ID) architecture to test optimizer effectiveness in isolation.
-    • Only SFT (RecLLM): The Native-LLM model trained only with Supervised Fine-Tuning.
-    • RecLLM: The proposed framework combining the Native-LLM architecture with the SAGE optimizer.
-4.1.3. Evaluation Metrics
-We evaluate performance across three dimensions using top-$K$ metrics (default $K=10$):
-  • Accuracy: Normalized Discounted Cumulative Gain (NDCG@K) and Recall (Recall@K).
-  • Diversity: Intra-List Diversity (ILD), measuring the average semantic distance between items within a generated slate, and Entropy@K, measuring the distribution of recommended categories across the test set.Since all items in a dataset category belong to the same root (e.g., ”Beauty”), calculating entropy at the root level is trivial.
-Therefore, we utilize the second-level sub-categories (e.g., ”Skin Care”, ”Makeup”, ”Fragrance” for Beauty) provided in the item metadata to calculate Shannon Entropy.
-This measures whether the model recommends a diverse range of granular interests rather than converging to a single sub-niche.
-  • Cold-Start: Given that precise item publication timestamps are unavailable in this dataset, we adopt the standard academic definition for cold-start evaluation.
-We define Cold Items as the set of items belonging to the bottom 20% of the interaction frequency distribution in the training set.
-This metric evaluates the model’s ability to discover and recommend long-tail items that have sparse historical interactions, preventing the ”rich-get-richer” feedback loop.
-4.1.4. Implementation Details
-The base model is Qwen3-8B.
-We set the slate length $L=6$.
-The SFT stage is trained for 1 epochs.
-In the RLHF stage, we use a learning rate of 5e65e-6 with the AdamW optimizer.
-For SAGE specific hyperparameters, based on validation set tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$, the Diversity Temperature $\beta=0.5$.
-4.2. Main Results
-Superiority of RecLLM: As shown in Table  1, our proposed method (”RecLLM”) consistently outperforms all baselines across all datasets.
-Specifically, compared to the strongest baseline OneRec-V2, RecLLM achieves improvements of 4.3% to 8.9% in NDCG metrics.
-This confirms that directly reusing the Native LLM vocabulary—when optimized correctly—can surpass models relying on heavy, separate tokenization.
-The Necessity of SAGE: Notably, the ”Only SFT” model performs poorly, significantly lagging behind even SASRec.
-SFT mainly functions to activate the model’s fundamental capabilities.
-without reinforcement learning to align the Native LLM with recommendation goals, the model fails to capture complex user preferences.
-Furthermore, OneRec-SAGE (applying SAGE to the old architecture) shows mixed results: while competitive, it slightly underperforms OneRec-V2 in some metrics (e.g., R@5 on Beauty).
-This suggests that SAGE’s aggressive exploration requires the robust semantic representation of the RecLLM to fully realize its potential without destabilizing the training.
-Table  2 investigates the impact of optimization strategies on recommendation diversity (Entropy) and long-tail item discovery (Cold-Recall).
-Breaking the Symmetric Conservatism: Comparing RecLLM against OneRec-V2, we observe substantial gains:
-  • Cold-Recall: SAGE improves cold-start performance by 49% (Beauty), 63% (Sports) and 43% (Toys).
-This validates that the Positive Boost Factor ($\epsilon_{boost}$) effectively allows high-potential fresh items to break through the symmetric gradient barrier of GBPO.
-  • Diversity: Entropy improves by 5.9% , 10.36% and 7.36%, confirming that the Entropy-Aware Penalty prevents the model from collapsing into safe, popular items.
-The Accuracy-Diversity Trade-off: It is worth noting that OneRec-SAGE achieves the highest absolute values for Entropy and Cold-Recall.
-However, revisiting Table  1, we see that OneRec-SAGE suffers a drop in top-k accuracy (e.g., lower R@5 than OneRec-V2).
-This indicates that applying aggressive exploration on the limited semantic ID vocabulary can lead to over-exploration.
-In contrast, RecLLM achieves the most Optimal state: it delivers the highest overall accuracy table 1 while still providing a massive upgrade in diversity and cold-start handling compared to the current SOTA baseline (OneRec-V2).
+with a fixed slate length $L=10$.
+Physical Meaning: this formulation smooths intra-item token fluctuations and can be interpreted as the average probability gain of generating the entire slate under the new policy relative to the old policy.
 Table 1. Overall performance comparison on three Amazon Product Reviews datasets. Bold indicates the best performance, and underline indicates the second best.
@@
 Method R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
-SASRec 0.0393 0.0639 0.0209 0.289 0.0240 0.0389 0.0130 0.0178 0.0420 0.0658 0.0217 0.0294
-ReaRec 0.0488 0.0702 0.0341 0.0231 0.0348 0.0152 0.0189 0.0249 0.0517 0.0706 0.0369 0.0430
+SASRec 0.0393 0.0639 0.0209 0.0289 0.0240 0.0389 0.0130 0.0178 0.0420 0.0658 0.0217 0.0294
+ReaRec 0.0488 0.0702 0.0341 0.0409 0.0231 0.0348 0.0152 0.0189 0.0517 0.0706 0.0369 0.0430
 TIGER 0.0413 0.0628 0.0277 0.0346 0.0216 0.0331 0.0145 0.0182 0.0367 0.0527 0.0255 0.0307
 LC-Rec 0.0495 0.0764 0.0338 0.0424 0.0269 0.0418 0.0177 0.0225 0.0350 0.0549 0.0221 0.0285
-OneRec-V2 0.0646 0.0929 0.0456 0.0550 0.0365 0.0551 0.0252 0.0317 0.0693 0.0958 0.0496 0.0581
-OneRec-SAGE 0.0612 0.0931 0.0424 0.0557 0.0351 0.0541 0.0246 0.0297 0.0671 0.0960 0.0466 0.0583
-Only SFT (RecLLM) 0.0291 0.023 0.0158 0.204 0.0163 0.0263 0.0104 0.0101 0.0386 0.0571 0.0169 0.0212
-RecLLM 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
-Improve +5.67% +4.7% +8.9% +7.2% +5.46% +4.33% +7.35% +6.96% +6.5% +6.3% +7.48% +7.11%
-Table 2. Relative to OneRec‑v2, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
+OneRec-GRPO 0.0646 0.0924 0.0456 0.0545 0.0365 0.0547 0.0252 0.0310 0.0693 0.0953 0.0496 0.0579
+OneRec-GBPO 0.0649 0.0928 0.0461 0.0550 0.0368 0.0551 0.0255 0.0317 0.0698 0.0958 0.0499 0.0581
+OneRec-SAGE 0.0651 0.0932 0.0463 0.0557 0.0371 0.0553 0.0256 0.0320 0.0701 0.0960 0.0502 0.0585
+Only SFT (TextRec) 0.0291 0.0495 0.0158 0.0204 0.0163 0.0263 0.0104 0.0139 0.0386 0.0571 0.0169 0.0212
+TextRec 0.0683 0.0973 0.0497 0.0590 0.0385 0.0575 0.0271 0.0339 0.0738 0.1018 0.0533 0.0622
+Improve +5.23% +4.84% +7.8% +7.27% +4.62% +4.35% +6.27% +6.94% +5.73% +6.26% +6.81% +7.05%
+Table 2. Relative to OneRec‑GBPO, SAGE delivers significant gains in cold‑start recovery and recommendation diversity.
 Dataset Beauty Sports Toys
-Method Entropy@10 Cold-Recall Entropy@10 Cold-Recall Entropy@10 Cold-Recall
-OneRec-V2 2.285 0.0326 2.274 0.0282 2.315 0.0365
-OneRec-SAGE 2.551 0.0619 2.528 0.0713 2.574 0.0695
-RecLLM 2.420 0.0487 2.510 0.0460 2.485 0.0522
-Improve +5.9% +49% +10.36% +63% +7.36% +43%
-4.3. Ablation Study
-To understand the contribution of each component, we conduct an ablation study on the Amazon-Beauty dataset. The results are visually presented in Figure  5.
+Method Entropy@10 Cold-Recall@10 Entropy@10 Cold-Recall@10 Entropy@10 Cold-Recall@10
+OneRec-GBPO 2.285 0.326 2.274 0.282 2.315 0.365
+OneRec-SAGE 2.551 0.619 2.528 0.567 2.574 0.705
+TextRec 2.420 0.487 2.510 0.460 2.485 0.522
+Improve +11.64% +89.9% +11.17% +101% +11.19% +93.2%
+Table 3. Performance on RecIF-Bench. Bold indicates the best and underline indicates the second best.
+Method Short Video Rec Ad Rec Product Rec Label Pred. Interactive Rec Label-Cond. Rec
+Pass@1 Pass@32 Recall@32 Pass@1 Pass@32 Recall@32 Pass@1 Pass@32 Recall@32 AUC Pass@1 Pass@32 Recall@32 Pass@32 Recall@32
+SASRec 0.0045 0.1003 0.0119 0.0044 0.0980 0.0293 0.0052 0.0914 0.0175 0.6244 - - - 0.0380 0.0140
+ReaRec 0.0052 0.1002 0.0120 0.0035 0.1054 0.0327 0.0030 0.0907 0.0189 0.6204 - - - 0.0381 0.0137
+TIGER 0.0168 0.1061 0.0132 0.0125 0.1769 0.0581 0.0120 0.1276 0.0283 0.6675 - - - 0.0337 0.0123
+LC-Rec 0.0341 0.1306 0.0180 0.0197 0.2096 0.0723 0.0178 0.1809 0.0416 0.6139 0.0890 0.3730 0.2394 0.0420 0.0170
+OneRec-GRPO 0.0548 0.2122 0.0369 0.0259 0.2700 0.0964 0.0223 0.2290 0.0538 0.6912 0.1250 0.5080 0.3458 0.0549 0.0235
+OneRec-GBPO 0.0551 0.2134 0.0377 0.0264 0.2731 0.0979 0.0227 0.2298 0.0545 0.6918 0.1256 0.5096 0.3481 0.0553 0.0248
+OneRec-SAGE 0.0554 0.2142 0.0381 0.0266 0.2733 0.0983 0.0229 0.2302 0.0554 0.6919 0.1258 0.5107 0.3495 0.0561 0.0254
+Only SFT (TextRec) 0.0049 0.0920 0.0107 0.0041 0.0940 0.0235 0.0032 0.0959 0.0161 0.6028 0.0410 0.2680 0.1660 0.0334 0.0128
+TextRec 0.0574 0.2196 0.0392 0.0273 0.2798 0.1015 0.0231 0.2410 0.0568 0.7017 0.1310 0.5280 0.3648 0.0574 0.0261
+4. Experiments
+We conduct experiments on both public academic benchmarks and a large-scale industrial benchmark to evaluate SAGE as a general-purpose optimizer for list-wise generative recommendation.
+Our goal is to verify that SAGE (i) improves accuracy while unlocking cold-start learning signals, (ii) mitigates diversity collapse under rejection-dominated feedback, and (iii) remains numerically stable.
+We evaluate SAGE across two representative action spaces: Semantic-ID generation (OneRec-style) and native-text generation (TextRec-style).
+4.1. Experimental Setup
+4.1.1. Datasets
+Amazon Product Reviews.
+We evaluate sequential recommendation on three real-world subsets from the Amazon Product Reviews dataset  (McAuley et al., 2015), which contains user reviews and item metadata from May 1996 to July 2014.
+We use the three categories Sports, Beauty, and Toys.
+Table 4. Diversity and cold-start evaluation on RecIF-Bench.
+Method Short Video
+Entropy@10 Cold-Recall@10
+OneRec-GBPO 4.284 0.521
+OneRec-SAGE 4.455 0.637
+TextRec 4.317 0.594
+RecIF-Bench.
+In addition, we evaluate on RecIF-Bench  (Zhou et al., 2025a), an instruction-following benchmark for generative recommendation foundation models.
+RecIF-Bench contains approximately 120M interactions from 200K users across three industrial domains: Short Video (impression sequences with multi-behavior feedback such as like/follow/comment/effective view/dislike), Ad (click sequences), and Product (click sequences).
+The benchmark provides (i) user portraits written as narrative text interleaved with behavioral evidence, (ii) item-side metadata and multimodal representations (text and visual embeddings; dense captions for a large portion of videos), and (iii) interaction-side labels enabling behavior-conditional modeling.
+We follow the official strict user-based split protocol: 20% users are held out for testing with zero user leakage; for each user, the interaction sequence is temporally split into history context and future targets.
+We report results on the core five recommendation tasks (Short Video Rec, Ad Rec, Product Rec, Interactive Rec, and Label-Conditional Rec) and the pointwise Label Prediction task.
+Native-text action space (selection within candidate pool).
+In TextRec, the policy does not freely generate arbitrary item strings.
+Instead, for each context $c$ we provide a candidate pool $\mathcal{I}(c)$ with textual item renderings (e.g., title/tags/category keywords), and the model outputs a length-$L$ slate by selecting items from $\mathcal{I}(c)$ (with deterministic parsing and validity checks).
+Outputs that do not match any candidate are treated as invalid and discarded (or mapped to a special null action), ensuring a well-defined and reproducible action space.
+4.1.2. Baselines
+To validate the effectiveness of our approach, we compare against three groups of competitive baselines, covering classical discriminative sequential recommenders, recent generative/retrieval-augmented recommenders, and OneRec-family variants.
+Discriminative sequential recommenders: SASRec (Kang and McAuley, 2018). These methods are inherently task-specific, and thus require training a separate model for each RecIF task.
+Generative / retrieval-augmented recommenders: ReaRec (Tang et al., 2025), TIGER (Rajput et al., 2023), and LC-Rec (Zheng et al., 2024).
+Following prior practice, we adapt these methods with task-specific modifications to support different RecIF tasks.
+For a fair comparison at the foundation-model scale, we implement LC-Rec as LC-Rec-8B by instantiating it with a comparable Qwen3-8B backbone.
+OneRec variants and our method: OneRec-GRPO (Zhou et al., 2025a) (Semantic IDs + GRPO), OneRec-GBPO (Zhou et al., 2025b) (Semantic IDs + GBPO), OneRec-SAGE (applying SAGE on the Semantic-ID architecture), Only SFT (TextRec) (native-vocabulary model trained only with supervised fine-tuning), and TextRec (native-vocabulary model further optimized with SAGE).
+4.1.3. Evaluation Metrics
+Amazon metrics.
+We evaluate top-$K$ performance (default $K=10$) from three perspectives: (i) Accuracy using NDCG@K and Recall@K; (ii) Diversity using Entropy@K.
+Since all items in a dataset category share the same root label (e.g., ”Beauty”), entropy at the root level is trivial; therefore, we compute Shannon Entropy using second-level sub-categories from item metadata (e.g., Skin Care, Makeup, Fragrance); (iii) Cold-start using Cold-Recall@K, where cold items are defined as the bottom 5% of the interaction-frequency distribution in the training set.
+RecIF-Bench metrics.
+We follow the official RecIF-Bench evaluation  (Zhou et al., 2025a).
+For recommendation tasks, we using Pass@K and Recall@K with $K\in\{1,32\}$.
+For Label Prediction, we using AUC.
+4.1.4. Implementation Details
+Backbone and generation. The base model is Qwen3-8B. We set slate length to $L=10$.
+Training recipe.
+We use a two-stage pipeline (SFT then RLHF) as described in Section 3.
+The SFT stage is trained for 1 epoch.
+In the RLHF stage, we use AdamW with a learning rate of 5e65e-6.
+For SAGE hyperparameters, based on validation tuning, we set the Positive Boost Factor $\epsilon_{boost}=0.3$ and Diversity Temperature $\beta=0.5$.
+RecIF-Bench specific settings.
+For TextRec, items are represented purely by native-text (title/tags/category keywords) without introducing a separate item-token vocabulary.
+For interactive and conditional recommendation, $L=10$, and we generate candidates using nucleus sampling with temperature 0.9.
+For OneRec-GRPO, OneRec-GBPO and OneRec-SAGE, we use the public checkpoints and evaluation settings in OpenOneRec  (Zhou et al., 2025a).
+For task-specific sequential baselines and TIGER, we reuse the reported benchmark numbers under the same protocol.
+4.2. Main Results
+4.2.1. Results on Amazon Product Reviews
+Table 1 shows that TextRec consistently achieves the best top-$K$ accuracy across Beauty, Sports, and Toys.
+Compared to the strongest baseline OneRec-GBPO, TextRec improves NDCG by 6.27% to 7.8%, indicating that directly reusing the native LLM vocabulary—when paired with preference optimization—can outperform architectures relying on separate Semantic-ID tokenizers.
+The Only SFT (TextRec) variant performs poorly and lags behind SASRec, suggesting that SFT mainly teaches formatting and coarse preference patterns; without preference optimization, the model fails to align generation with ranking objectives.
+Applying SAGE to the Semantic-ID architecture (OneRec-SAGE) yields competitive but mixed accuracy changes, implying that aggressive exploration is harder to stabilize when the output space is constrained by a discrete Semantic-ID vocabulary.
+Beyond accuracy, Table 2 highlights SAGE’s effect on long-tail discovery and diversity.
+Relative to OneRec-GBPO, SAGE yields large gains in Cold-Recall and also improves Entropy, supporting our diagnosis that GBPO-style symmetric clipping tends to suppress positive learning signals for rare items and encourages conservative, popularity-biased updates.
+4.2.2. Results on RecIF-Bench
+Table 3 reports results on six RecIF-Bench tasks.
+TextRec consistently outperforms OneRec-GBPO across the core recommendation tasks, with improvements that are especially clear on instruction-conditioned settings (e.g., Interactive Rec and Label-Conditional Rec).
+This indicates that representing items with native-text can better exploit the backbone LLM’s semantic priors when the task is framed as instruction following rather than pure next-item prediction.
+The Only SFT variant again underperforms, consistent with the Amazon findings: SFT alone does not provide sufficiently strong preference signals to drive high-quality ranking under the RecIF-Bench protocol.
+OneRec-SAGE remains competitive but does not consistently dominate top-$K$ accuracy, reinforcing the empirical pattern that exploration can become over-aggressive when the action space is limited by a Semantic-ID vocabulary.
+Finally, Table 4 evaluates diversity and cold-start on RecIF-Bench (Short Video).
+Relative to OneRec-GBPO, TextRec improves both Entropy@10 and Cold-Recall@10, while OneRec-SAGE achieves the highest absolute values on these two metrics.
+Taken together with the accuracy table, this suggests a consistent trade-off: SAGE increases exploration and long-tail coverage; the native-vocabulary TextRec setting better preserves accuracy under such exploration than Semantic-ID generation.
 Figure 5.
@@
 Removing individual components leads to degradation in their respective targeted areas (e.g., removing Boost hurts Cold-Start most), validating the design of each module.
-  • w/o Positive Boost: Setting $\epsilon_{boost}=0$ reverts the positive boundary to standard GBPO ($r\leq 1$).
-We observe a drastic drop in Cold-Start Recall (Green Bar) to 0.50 of the full model’s performance, confirming this component’s critical role in item exploration.
-  • w/o Entropy Penalty: Removing the entropy term and using the static GBPO boundary for negative feedback leads to a significant decrease in Entropy@10 (Orange Bar) to 0.78, indicating a collapse towards redundant recommendations.
-  • w/o Signal Decoupling: Using a simple weighted sum of rewards instead of decoupled normalization results in a noticeable drop in NDCG@10 (Blue Bar) to 0.90 and a decrease in other metrics.
-This suggests that ”Reward Collapse” was indeed hindering the precise multi-objective optimization.
-The full SAGE model achieves the best balance, demonstrating that the asymmetric adaptive dynamics and signal decoupling are complementary and essential.
+4.3. Ablation Study
+To isolate the contribution of each component, we conduct ablations on Amazon-Beauty. Figure 5 normalizes metrics to the full model.
+Removing Positive Boost ($\epsilon_{boost}=0$) drastically reduces Cold-Start Recall, showing that relaxing the conservative positive update boundary is crucial for long-tail discovery.
+Removing the Entropy Penalty substantially decreases Entropy@10, indicating that diversity-aware negative feedback prevents collapse into redundant popular outputs.
+Replacing Signal Decoupling with a simple weighted reward sum reduces NDCG@10 and other metrics, supporting the hypothesis that multi-objective reward mixing can cause signal interference and unstable optimization.
+Overall, the best results require all three components working together.
 5. Conclusion
-In this paper, we identified fundamental limitations in existing optimization strategies for generative recommendation, specifically the ”Symmetric Conservatism” of Gradient-Bounded Policy Optimization (GBPO), which leads to cold-start suppression and diversity collapse.
-To address these challenges within a list-wise generation paradigm, we proposed SAGE (Sequence-level Adaptive Gradient Evolution).
-SAGE introduces a novel theoretical framework featuring **Asymmetric Adaptive Dynamics**.
-By applying a ”Positive Boost” to high-potential successful slates and an ”Entropy-Aware Penalty” to low-diversity failures, SAGE dynamically aligns the optimization trajectory with the long-tailed, high-noise nature of recommendation feedback.
-Furthermore, we incorporated Multi-Objective Signal Decoupling to ensure high-fidelity gradient signals in complex real-world scenarios.
-RecLLM demonstrates that native‑vocabulary reuse not only surpasses models dependent on large standalone tokenizers but also maximizes the utility of the underlying LLM.
-While the absence of an independent vocabulary necessitates language‑level item encoding—posing challenges for real‑time systems—this constraint is expected to fade with increasing computational capacity.
+We studied optimization failures in generative recommendation, focusing on the ”Symmetric Conservatism” induced by GBPO-style bounded updates, which empirically manifests as cold-start suppression and diversity collapse.
+To address this in list-wise generation, we proposed SAGE (Sequence-level Adaptive Gradient Evolution), featuring (i) Asymmetric Adaptive Dynamics via a Positive Boost for high-potential successful slates and an Entropy-Aware Penalty for low-diversity failures, and (ii) Multi-Objective Signal Decoupling to preserve gradient fidelity under mixed feedback.
+Across three Amazon Product Reviews datasets and the large-scale industrial RecIF-Bench benchmark, TextRec—which reuses the LLM’s native vocabulary and is optimized with SAGE—consistently achieves the strongest overall accuracy while substantially improving long-tail coverage and recommendation diversity relative to OneRec-GBPO.
+The results also reveal a practical stability pattern: applying strong exploration to a constrained Semantic-ID action space can increase diversity but may sacrifice top-$K$ accuracy, whereas native-text generation better maintains the accuracy–diversity balance under SAGE.
+A remaining limitation is that native-vocabulary recommendation requires language-level item rendering, which can be challenging for latency-sensitive systems; nevertheless, the consistent gains on both academic and industrial benchmarks suggest that native-text generation is a promising direction as compute and deployment tooling improve.